OpenStudy (anonymous):

It is known that making one 1lb of honey requires 20,000 trips by bees to flowers to gather nectar. How many pounds of honey would 85,000 trips produce?

OpenStudy (anonymous):

simply divide 85,000 by 20,000 to find the amount if I'm not mistaken

OpenStudy (anonymous):

ratio 1/20000 = x / 85000 times 85000 both sides 85000/20000 = x x = 4.25 4.25 pounds

OpenStudy (anonymous):

Think about a simpler problem. How much honey would 40,000 trips produce? What about 60,000 trips? What about 100,000 trips? You see a pattern? How much honey would 20,000 * k trips produce, if k is some random number? So how much honey would 85,000 trips produce? If you could solve the previous problem, then you’ll see the solution to the problem you asked.

OpenStudy (anonymous):

thanks everyone for the tips and answers